Define a relation R on the set of integers Z
by R= {(n, n+ 3k) I k belongs to Z}. Show that R is an equivalence relation. Also find all distinct
equivalence classes.
1
Expert's answer
2020-06-10T19:48:00-0400

By given relation, "aRb" if "b = a+3k, k \in \Z" .

Reflexive: "a = a+3 (0)" "\implies aRa" . So "R" is an reflexive relation.

Symmetric: Let "aRb \implies b = a+3k" where "k" is an integer.

"\implies a = b - 3k = b + 3(-k) = b + 3k_1" and "k_1 = -k \in \Z" "\implies bRa".

So. "R" is an symmetric relation.

Transitive: Let "aRb, bRc" "\implies b = a+3k_1 , c =b +3k_2"

"\implies c = a+3k_1 +3k_2 = a+3(k_1+k_2) = a+3k" where "k = k_1+k_2 \in \Z".

Hence, "aRc" . So, "R" is an transitive relation.

Thus, "R" is an equivalence relation.


Equivalence class of 0 = "\{3k: k\in \Z\} = \{ \_ \ \_ \ \_, -6, -3, 0, 3 , 6, \_ \ \_ \ \_ \}"

Equivalence class of 1 = "\{3k+1 : k\in \Z\} = \{ \_ \ \_ \ \_, -5, -2, 1 , 4 , 7, \_ \ \_ \ \_ \}"

Equivalence class of 2 = "\{3k+2 : k\in \Z\} = \{ \_ \ \_ \ \_, -4, -1, 2 , 5 , 8, \_ \ \_ \ \_ \}"

These three distinct equivalence classes are possible.